# Performance Analysis of Physical Layer Security: From Far-Field to Near-Field

### v1

@@ III-B2 USW Model
for $i\in\left\{x,z\right\}$, and where $b_{x}=\frac{2\pi d}{\lambda}\left(\Phi_{\mathsf{b}}-\Phi_{\mathsf{e}}\right)$, $b_{z}=\frac{2\pi d}{\lambda}\left(\Omega_{\mathsf{b}}-\Omega_{\mathsf{e}}\right)$, $a_{x}=\frac{\pi d^{2}}{\lambda}\left(\frac{1-\Phi_{\mathsf{b}}^{2}}{r_{\mathsf{b}}}-\frac{1-\Phi_{\mathsf{e}}^{2}}{r_{\mathsf{e}}}\right)$, $a_{z}=\frac{\pi d^{2}}{\lambda}\left(\frac{1-\Omega_{\mathsf{b}}^{2}}{r_{\mathsf{b}}}-\frac{1-\Omega_{\mathsf{e}}^{2}}{r_{\mathsf{e}}}\right)$, and $\mathrm{erf}\left(x\right)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}{\mathrm{e}^{-t^{2}}\mathrm{d}t}$ is the error function [22, Eq. (8.250.1)]. The expression for the secrecy capacity is shown in Table I.
@@ III-B3 NUSW Model
Due to the linearly varying phase shifts in the UPW model, the channels for Bob and Eve become parallel when they are located in the same direction, resulting in $\rho=1$ and thus a secrecy capacity ${\mathsf{C}}=0$. In contrast, the nonlinear varying phase shifts in the spherical-wave models, i.e., USW and NUSW, prevent the channels for Bob and Eve from aligning perfectly unless they are at the exact same location. According to Corollary 3, we draw the following conclusion.
@@
$\lim_{M_{x},M_{z}\rightarrow\infty}G_{k}^{\mathsf{N}}=\frac{4A\lim\limits_{x,z\rightarrow\infty}\arctan\left(\frac{xz\Psi_{k}^{-1}}{\sqrt{\Psi_{k}^{2}\!+\!x^{2}\!+\!z^{2}}}\right)}{4\pi d^{2}}=\frac{A}{2d^{2}}.$
@@
For comparison, we summarize skey outcomes from the above analysis across different channel models in Table II, leading to the following observations.
@@ III-C Depth of Insecurity
$\!\!\mathbf{Q}_{\mathsf{e}}^{-1}\mathbf{h}_{\mathsf{b}}$ $=\left(\mathbf{I}+\overline{\gamma}_{\mathsf{e}}\mathbf{h}_{\mathsf{e}}\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}\right)^{-1}\mathbf{h}_{\mathsf{b}}$ (45a)
=$\!\!\overset{\left(a\right)}{=}\!\bigg{(}\mathbf{I}\!-\frac{\overline{\gamma}_{\mathsf{e}}\mathbf{h}_{\mathsf{e}}\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}}{1+\overline{\gamma}_{\mathsf{e}}\!\left\|\mathbf{h}_{\mathsf{e}}\right\|^{2}}\bigg{)}\mathbf{h}_{\mathsf{b}}\overset{\left(b\right)}{\approx}\!\bigg{(}\mathbf{I}-\frac{\mathbf{h}_{\mathsf{e}}\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}}{\left\|\mathbf{h}_{\mathsf{e}}\right\|^{2}}\bigg{)}\mathbf{h}_{\mathsf{b}},$ (45b)
@@
$\mathbf{w}_{\infty}^{\star}$ $=\sqrt{P}\eta(\mathbf{I}-{\mathbf{h}_{\mathsf{e}}\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}}/{\|\mathbf{h}_{\mathsf{e}}\|^{2}})\mathbf{h}_{\mathsf{b}}$ (46a)
$=\sqrt{P}\eta(\mathbf{h}_{\mathsf{b}}-({\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}\mathbf{h}_{\mathsf{b}}}/{\|\mathbf{h}_{\mathsf{e}}\|^{2}})\mathbf{h}_{\mathsf{e}}),$ (46b)
where $\eta=\|(\mathbf{I}-{\mathbf{h}_{\mathsf{e}}\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}}/{\|\mathbf{h}_{\mathsf{e}}\|^{2}})\mathbf{h}_{\mathsf{b}}\|^{-1}$ is for normalization. Notably, $\mathbf{w}_{\infty}^{\star}$ is orthogonal to $\mathbf{h}_{\mathsf{e}}$, i.e.,
$\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}\mathbf{w}_{\infty}^{\star}=\sqrt{P}\eta\left(\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}\mathbf{h}_{\mathsf{b}}-{\|\mathbf{h}_{\mathsf{e}}\|^{2}}/{\|\mathbf{h}_{\mathsf{e}}\|^{2}}\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}\mathbf{h}_{\mathsf{b}}\right)=0.$ (47)
@@
The results in (46b) suggest that in the high-SNR regime, the optimal beamformer aligns with $\mathbf{h}_{\mathsf{b}}$ minus its projection on $\mathbf{h}_{\mathsf{e}}$, which represents the vector most aligned with $\mathbf{h}_{\mathsf{b}}$ within the null space of $\mathbf{h}_{\mathsf{e}}$.
@@
$\cos\psi=\frac{\left|\|\mathbf{h}_{\mathsf{b}}\|-\frac{|\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}\mathbf{h}_{\mathsf{b}}|^{2}}{\|\mathbf{h}_{\mathsf{b}}\|\|\mathbf{h}_{\mathsf{e}}\|^{2}}\right|^{2}}{\mathbf{h}_{\mathsf{b}}^{\mathsf{H}}\left(\mathbf{I}-\frac{\mathbf{h}_{\mathsf{e}}\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}}{\left\|\mathbf{h}_{\mathsf{e}}\right\|^{2}}\right)\mathbf{h}_{\mathsf{b}}}=\frac{\|\mathbf{h}_{\mathsf{b}}\|^{2}\left|1-\frac{|\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}\mathbf{h}_{\mathsf{b}}|^{2}}{\|\mathbf{h}_{\mathsf{b}}\|^{2}\|\mathbf{h}_{\mathsf{e}}\|^{2}}\right|^{2}}{\|\mathbf{h}_{\mathsf{b}}\|^{2}\left(1-\frac{|\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}\mathbf{h}_{\mathsf{b}}|^{2}}{\|\mathbf{h}_{\mathsf{b}}\|^{2}\|\mathbf{h}_{\mathsf{e}}\|^{2}}\right)}.$
@@
Please refer to Appendix -E for more details. ∎
@@ IV-A Minimum Power Requirement Characterization
$\min_{{\mathbf{v}},t}~{}t~{}~{}{\rm{s.t.}}~{}\frac{1\!+t\sigma_{\mathsf{b}}^{-2}\lvert{\mathbf{v}}^{\mathsf{H}}{\mathbf{h}}_{\mathsf{b}}\rvert^{2}}{1\!+t\sigma_{\mathsf{e}}^{-2}\lvert{\mathbf{v}}^{\mathsf{H}}{\mathbf{h}}_{\mathsf{e}}\rvert^{2}}\geq 2^{{\mathsf{R}_{0}}},~{}t>0,~{}\lVert{\mathbf{v}}\rVert^{2}\!=\!1.$ (51)
Multiplying both sides of the first constraint in (51) by $1+t\sigma_{\mathsf{e}}^{-2}\lvert{\mathbf{v}}^{\mathsf{H}}{\mathbf{h}}_{\mathsf{e}}\rvert^{2}$ gives
@@
where $\mu_{\bm{\Theta}}$ denotes the principal eigenvalue of ${\bm{\Theta}}$. Upon combining (52) and (53), we find that $t$ must satisfy $t\geq\frac{2^{{\mathsf{R}_{0}}}-1}{\mu_{\bm{\Theta}}}>0$ to guarantee the feasibility of problem (51). This means that the minimum value of $t$, i.e., the minimum required power, can be written as follows:
$\mathsf{P}=({2^{{\mathsf{R}_{0}}}-1})/{\mu_{\bm{\Theta}}}.$ (54)
@@
$\mathsf{P}={2\left(2^{\mathsf{R}_{0}}-1\right)}/({\xi+\sqrt{\xi^{2}+\chi}}),$ (55)
@@
Please refer to Appendix -F for more details. ∎
@@
Since the minimum required power $\mathsf{P}$ is also expressed as a function of the channel gains and the channel correlation factor, which were derived in the previous section, the subsequent text will focus solely on the asymptotic analysis as $M_{x},M_{z}\rightarrow\infty$ ($M\rightarrow\infty$) for the various far-field and near-field channel models. For brevity, we assume $\sigma_{\mathsf{b}}^{2}=\sigma_{\mathsf{e}}^{2}=\sigma^{2}$.

### v2

@@ III-B1 UPW Model
Figure 3: Illustration of the ULA.
Next, we consider a special case where the BS is equipped with a uniform linear array (ULA). In this case, we have $M_{x}=1$ and $M_{z}=M$, as illustrated in Fig. 3.
Corollary 7.
With a ULA, the channel gains under the UPW model are given by $G_{k}=\frac{MA\Psi_{k}}{4\pi r_{k}^{2}}$ for $k\in\{\mathsf{b},\mathsf{e}\}$, and the channel correlation factor is given by
$\rho=\begin{cases}1&\phi_{\mathsf{b}}=\phi_{\mathsf{e}}\\ \frac{1-\cos\left(M\varDelta_{\Omega}\right)}{M^{2}\left(1-\cos\varDelta_{\Omega}\right)}&\phi_{\mathsf{b}}\neq\phi_{\mathsf{e}}\\ \end{cases}$ (34)
Proof:
Similar to the proof of Theorem 2. ∎
Having derived the expressions of the channel gains and the correlation factor for the ULA case, we can readily obtain the secrecy capacity and perform the similar analysis as in the UPA scenario, which are omitted due to space limitation.
@@ III-B2 USW Model
for $i\in\left\{x,z\right\}$, and where $b_{x}=\frac{2\pi d}{\lambda}\left(\Phi_{\mathsf{b}}-\Phi_{\mathsf{e}}\right)$, $b_{z}=\frac{2\pi d}{\lambda}\left(\Omega_{\mathsf{b}}-\Omega_{\mathsf{e}}\right)$, $a_{x}=\frac{\pi d^{2}}{\lambda}\left(\frac{1-\Phi_{\mathsf{b}}^{2}}{r_{\mathsf{b}}}-\frac{1-\Phi_{\mathsf{e}}^{2}}{r_{\mathsf{e}}}\right)$, $a_{z}=\frac{\pi d^{2}}{\lambda}\left(\frac{1-\Omega_{\mathsf{b}}^{2}}{r_{\mathsf{b}}}-\frac{1-\Omega_{\mathsf{e}}^{2}}{r_{\mathsf{e}}}\right)$, and $\mathrm{erf}\left(x\right)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}{\mathrm{e}^{-t^{2}}\mathrm{d}t}$ is the error function [32, Eq. (8.250.1)]. The expression for the secrecy capacity is shown in Table I.
@@
We then consider the ULA case under the USW model.
Corollary 10.
When a ULA is employed, under the USW channel model, the channel gains are given by $G_{k}=\frac{MA\Psi_{k}}{4\pi r_{k}^{2}}$ for $k\in\{\mathsf{b},\mathsf{e}\}$, and the channel correlation factor satisfies $\rho=\frac{\delta_{z}}{M^{2}}$.
Proof:
Similar to the proof of Theorem 3. ∎
@@ III-B3 NUSW Model
Due to the linearly varying phase shifts in the UPW model, the channels for Bob and Eve become parallel when they are located in the same direction, resulting in $\rho=1$ and thus $\mathsf{C}=0$ for $r_{\mathsf{b}}\geq r_{\mathsf{e}}$. In contrast, the nonlinear varying phase shifts in the spherical-wave models, i.e., USW and NUSW, prevent the channels for Bob and Eve from aligning perfectly unless they are at the exact same location. According to Corollary 3, we draw the following conclusion.
@@
$\!\lim_{M_{x},M_{z}\rightarrow\infty}\!G_{k}^{\mathsf{N}}\!=\!\frac{4A\!\lim\limits_{x,z\rightarrow\infty}\!\arctan\!\Big{(}\frac{xz\Psi_{k}^{-1}}{\sqrt{\Psi_{k}^{2}\!+x^{2}\!+z^{2}}}\Big{)}}{4\pi d^{2}}\!=\frac{A}{2d^{2}}.$ (45)
@@
We next investigate the ULA case in the following corollary.
Corollary 13.
Under the NUSW model, the channel gain of node $k\in\{{\mathsf{b}},{\mathsf{e}}\}$ is given by
$\begin{split}G_{k}=&\frac{A\varepsilon_{k}\sin\phi_{k}}{4\pi d^{2}\sin\theta_{k}}\bigg{(}\frac{M\varepsilon_{k}-2\cos\theta_{k}}{\sqrt{M^{2}\varepsilon_{k}^{2}-4M\cos\theta_{k}\varepsilon_{k}+4}}\\ &+\frac{M\varepsilon_{k}+2\cos\theta_{k}}{\sqrt{M^{2}\varepsilon_{k}^{2}+4M\cos\theta_{k}\varepsilon_{k}+4}}\bigg{)}.\end{split}$ (47)
The channel correlation factor for the ULA case can be written as follows:
$\begin{split}&\rho=\frac{M^{2}A^{2}\Psi_{\mathsf{b}}\Psi_{\mathsf{e}}\pi^{2}}{16G_{\mathsf{b}}^{\mathsf{N}}G_{\mathsf{e}}^{\mathsf{N}}r_{\mathsf{b}}^{2}r_{\mathsf{e}}^{2}T^{4}}\\ &~{}~{}\times\left|\sum\nolimits_{t=1}^{T}{\sqrt{\left(1-\zeta_{t}^{2}\right)}}g^{\prime}_{1}\left(M\varepsilon_{\mathsf{b}}\zeta_{t}\right)g^{\prime}_{2}\left(M\varepsilon_{\mathsf{b}}\zeta_{t}\right)\right|^{2},\end{split}$ (48)
where $g^{\prime}_{i}(z)\triangleq g_{i}(0,z)$ for $i=1,2$.
Proof:
Please refer to Appendix -E for more details. ∎
For comparison, we summarize key outcomes from the above analysis across different channel models in Table II, leading to the following observations.
@@ III-C Depth of Insecurity
$\!\!\mathbf{Q}_{\mathsf{e}}^{-1}\mathbf{h}_{\mathsf{b}}$ $=\left(\mathbf{I}+\overline{\gamma}_{\mathsf{e}}\mathbf{h}_{\mathsf{e}}\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}\right)^{-1}\mathbf{h}_{\mathsf{b}}$
=$\overset{\left(a\right)}{=}\bigg{(}\mathbf{I}\!-\frac{\overline{\gamma}_{\mathsf{e}}\mathbf{h}_{\mathsf{e}}\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}}{1+\overline{\gamma}_{\mathsf{e}}\!\left\|\mathbf{h}_{\mathsf{e}}\right\|^{2}}\bigg{)}\mathbf{h}_{\mathsf{b}}$
$\overset{\left(b\right)}{\approx}\bigg{(}\mathbf{I}-\frac{\mathbf{h}_{\mathsf{e}}\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}}{\left\|\mathbf{h}_{\mathsf{e}}\right\|^{2}}\bigg{)}\mathbf{h}_{\mathsf{b}},$ (50)
@@
$\mathbf{w}_{\infty}^{\star}$ $=\sqrt{P}\eta\left(\mathbf{I}-\frac{\mathbf{h}_{\mathsf{e}}\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}}{\|\mathbf{h}_{\mathsf{e}}\|^{2}}\right)\mathbf{h}_{\mathsf{b}}=\sqrt{P}\eta\left(\mathbf{h}_{\mathsf{b}}-\frac{\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}\mathbf{h}_{\mathsf{b}}}{\|\mathbf{h}_{\mathsf{e}}\|^{2}}\mathbf{h}_{\mathsf{e}}\right),$ (51)
where $\eta=\|(\mathbf{I}-\frac{\mathbf{h}_{\mathsf{e}}\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}}{\|\mathbf{h}_{\mathsf{e}}\|^{2}})\mathbf{h}_{\mathsf{b}}\|^{-1}$ is for normalization. Notably, $\mathbf{w}_{\infty}^{\star}$ is orthogonal to $\mathbf{h}_{\mathsf{e}}$, i.e.,
$\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}\mathbf{w}_{\infty}^{\star}=\sqrt{P}\eta\left(\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}\mathbf{h}_{\mathsf{b}}-\frac{\|\mathbf{h}_{\mathsf{e}}\|^{2}}{\|\mathbf{h}_{\mathsf{e}}\|^{2}}\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}\mathbf{h}_{\mathsf{b}}\right)=0.$ (52)
@@
The results in (51) suggest that in the high-SNR regime, the optimal beamformer aligns with $\mathbf{h}_{\mathsf{b}}$ minus its projection on $\mathbf{h}_{\mathsf{e}}$, which represents the vector most aligned with $\mathbf{h}_{\mathsf{b}}$ within the null space of $\mathbf{h}_{\mathsf{e}}$.
@@
$\cos\psi=\frac{\left|\|\mathbf{h}_{\mathsf{b}}\|-\frac{|\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}\mathbf{h}_{\mathsf{b}}|^{2}}{\|\mathbf{h}_{\mathsf{b}}\|\|\mathbf{h}_{\mathsf{e}}\|^{2}}\right|^{2}}{\mathbf{h}_{\mathsf{b}}^{\mathsf{H}}\left(\mathbf{I}-\frac{\mathbf{h}_{\mathsf{e}}\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}}{\left\|\mathbf{h}_{\mathsf{e}}\right\|^{2}}\right)\mathbf{h}_{\mathsf{b}}}=\frac{\|\mathbf{h}_{\mathsf{b}}\|^{2}\left|1-\frac{|\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}\mathbf{h}_{\mathsf{b}}|^{2}}{\|\mathbf{h}_{\mathsf{b}}\|^{2}\|\mathbf{h}_{\mathsf{e}}\|^{2}}\right|^{2}}{\|\mathbf{h}_{\mathsf{b}}\|^{2}\left(1-\frac{|\mathbf{h}_{\mathsf{e}}^{\mathsf{H}}\mathbf{h}_{\mathsf{b}}|^{2}}{\|\mathbf{h}_{\mathsf{b}}\|^{2}\|\mathbf{h}_{\mathsf{e}}\|^{2}}\right)}.$ (54)
@@
Please refer to Appendix -F for more details. ∎
@@ IV-A Minimum Power Requirement Characterization
$\min_{{\mathbf{v}},t}~{}~{}$ $t$ (57a)
${\rm{s.t.}}~{}~{}$ $\frac{1+t\sigma_{\mathsf{b}}^{-2}\lvert{\mathbf{v}}^{\mathsf{H}}{\mathbf{h}}_{\mathsf{b}}\rvert^{2}}{1+t\sigma_{\mathsf{e}}^{-2}\lvert{\mathbf{v}}^{\mathsf{H}}{\mathbf{h}}_{\mathsf{e}}\rvert^{2}}\geq 2^{{\mathsf{R}_{0}}},~{}t>0,~{}\lVert{\mathbf{v}}\rVert^{2}=1.$ (57b)
Multiplying both sides of the first constraint in (57a) by $1+t\sigma_{\mathsf{e}}^{-2}\lvert{\mathbf{v}}^{\mathsf{H}}{\mathbf{h}}_{\mathsf{e}}\rvert^{2}$ gives
@@
where $\mu_{\bm{\Theta}}$ denotes the principal eigenvalue of ${\bm{\Theta}}$. Upon combining (58) and (59), we find that $t$ must satisfy $t\geq\frac{2^{{\mathsf{R}_{0}}}-1}{\mu_{\bm{\Theta}}}>0$ to guarantee the feasibility of problem (57a). This means that the minimum value of $t$, i.e., the minimum required power, can be written as follows:
$\mathsf{P}=\frac{2^{{\mathsf{R}_{0}}}-1}{\mu_{\bm{\Theta}}}.$ (60)
@@
$\mathsf{P}=\frac{2\left(2^{\mathsf{R}_{0}}-1\right)}{\xi+\sqrt{\xi^{2}+\chi}},$ (61)
@@
Please refer to Appendix -G for more details. ∎
@@
Based on the previous results and the above corollary, we have the following observation.
Remark 15.
Unlike FFC, NFC can achieve any desired secrecy rate for Bob, regardless of his location, provided there is sufficient transmit power.
Model Minimum Required Power
UPW $\frac{2\left(2^{\mathsf{R}_{0}}-1\right)}{\frac{MA\Psi_{\mathsf{b}}}{4\pi r_{\mathsf{b}}^{2}\sigma_{\mathsf{b}}^{2}}-\frac{2^{\mathsf{R}_{0}}MA\Psi_{\mathsf{e}}}{4\pi r_{\mathsf{e}}^{2}\sigma_{\mathsf{e}}^{2}}+\sqrt{\left(\frac{MA\Psi_{\mathsf{b}}}{4\pi r_{\mathsf{b}}^{2}\sigma_{\mathsf{b}}^{2}}-\frac{2^{\mathsf{R}_{0}}MA\Psi_{\mathsf{e}}}{4\pi r_{\mathsf{e}}^{2}\sigma_{\mathsf{e}}^{2}}\right)^{2}+\frac{2^{\mathsf{R}_{0}+2}M^{2}A^{2}\Psi_{\mathsf{b}}\Psi_{\mathsf{e}}}{\left(4\pi\right)^{2}r_{\mathsf{b}}^{2}r_{\mathsf{e}}^{2}\sigma_{\mathsf{b}}^{2}\sigma_{\mathsf{e}}^{2}}\left(1-\rho_{{}_{\mathsf{P}}}\right)}}$
USW $\frac{2\left(2^{\mathsf{R}_{0}}-1\right)}{\frac{MA\Psi_{\mathsf{b}}}{4\pi r_{\mathsf{b}}^{2}\sigma_{\mathsf{b}}^{2}}-\frac{2^{\mathsf{R}_{0}}MA\Psi_{\mathsf{e}}}{4\pi r_{\mathsf{e}}^{2}\sigma_{\mathsf{e}}^{2}}+\sqrt{\left(\frac{MA\Psi_{\mathsf{b}}}{4\pi r_{\mathsf{b}}^{2}\sigma_{\mathsf{b}}^{2}}-\frac{2^{\mathsf{R}_{0}}MA\Psi_{\mathsf{e}}}{4\pi r_{\mathsf{e}}^{2}\sigma_{\mathsf{e}}^{2}}\right)^{2}+\frac{2^{\mathsf{R}_{0}+2}M^{2}A^{2}\Psi_{\mathsf{b}}\Psi_{\mathsf{e}}}{\left(4\pi\right)^{2}r_{\mathsf{b}}^{2}r_{\mathsf{e}}^{2}\sigma_{\mathsf{b}}^{2}\sigma_{\mathsf{e}}^{2}}\left(1-\rho_{{}_{\mathsf{S}}}\right)}}$
NUSW $\frac{2\left(2^{\mathsf{R}_{0}}-1\right)}{\sigma_{\mathsf{b}}^{-2}G_{\mathsf{b}}^{\mathsf{N}}-2^{\mathsf{R}_{0}}\sigma_{\mathsf{e}}^{-2}G_{\mathsf{e}}^{\mathsf{N}}+\sqrt{\left(\sigma_{\mathsf{b}}^{-2}G_{\mathsf{b}}^{\mathsf{N}}-2^{\mathsf{R}_{0}}\sigma_{\mathsf{e}}^{-2}G_{\mathsf{e}}^{\mathsf{N}}\right)^{2}+2^{\mathsf{R}_{0}+2}\sigma_{\mathsf{b}}^{-2}\sigma_{\mathsf{e}}^{-2}G_{\mathsf{b}}^{\mathsf{N}}G_{\mathsf{e}}^{\mathsf{N}}\left(1-\rho_{{}_{\mathsf{N}}}\right)}}$
TABLE III: Minimum required transmission power under different channel models.
Note that minimum required power $\mathsf{P}$ is also expressed as a function of the channel gains and the channel correlation factor, which were derived in the previous section. We conclude the expressions of the minimum required power for different channel models in Table III.
In the following subsection, we focus on the asymptotic analysis as $M_{x},M_{z}\rightarrow\infty$ ($M\rightarrow\infty$) for the various far-field and near-field channel models to unveil further insights. For brevity, we assume $\sigma_{\mathsf{b}}^{2}=\sigma_{\mathsf{e}}^{2}=\sigma^{2}$.